: The students take a survey which has a single question with three possible answers; A, B, and C Suppose we have a sam ple of 200 esponses which provides 80 A, 70 B, and 50 C. Show the frequency and relative frequency distributions: Class Frequency Relative Frequency Total 2. A partial relative frequency distribution is given. Frequency % Frequency Class 36% 24% 28% D What is the relative frequency of class D? Assume the total sample size is 200. What is the frequency of class D? D Show the frequency distribution for each class.
Answer #1

(1)

Class Frequency Relative Frequency
A 80 80/200 = 0.40
B 70 70/200 = 0.35
C 50 50/200 = 0.25
Total 200 1.00

(2)

Class % Frequency Frequency
A 36 % 200 X 36%= 72
B 24 % 200 X 24% = 48
C 28 % 200 X 28% = 56
D 12% 200 X 12% = 24
Total 100 % 200

(1) Relative frequency of class D = 100% - (36% + 24% + 28% ) = 100% - 88% = 12%

(2) Frequency of class D = 200 X12% = 24

(3) Frequency distribution of each class is given in the above Table.